# Regarding the extension of metaplectic geometrical optics to modelling evanescent waves in ray-tracing codes

### v1

@@ Abstract
Metaplectic geometrical optics (MGO) is a recently developed ray-tracing framework to accurately compute the wavefield behavior near a caustic (turning point or focal point), where traditional ray-tracing breaks down. However, MGO has thus far been restricted to having real-valued wavevectors. This is disadvantageous because often upon crossing a caustic from the ‘illuminated’ region to the ‘shadow’ region, two real-valued rays coalesce into one complex-valued ray corresponding to the transition from propagating to evanescent behavior; MGO is currently unable to model such a transition. One can distinguish caustics as having either ‘illuminated shadows’ or ‘proper shadows’ – the former corresponds to when the caustic shadow still contains real-valued rays (albeit in a fewer quantity than in the illuminated region), while the latter corresponds to when the caustic shadow contains no real-valued rays. Here, by means of examples, we show how MGO can be used to model both types of caustic shadows. First, for illuminated shadows we show that MGO can actually be used ‘as is’, provided a corrected quadrature-angle bias is used in the numerical implementation compared to that proposed in the original references. This is then implemented and demonstrated in a recently developed MGO ray-tracing code. Second, we show that for proper shadows, the MGO formalism can still be used provided that the symplectic rotation matrix used to remove caustics along the rays is allowed to be complex-valued. In both cases, strong agreement is seen between the MGO solution and the exact solution, demonstrating the potential of MGO for improving the predictive capability of ray-tracing codes, and laying the foundations for modeling more complicated evanescent phenomena such as tunnelling with MGO.
@@ Acknowledgements
The work presented here is supported by the Carlsberg Foundation, grant CF23-0181.

### v2

@@ Abstract
Metaplectic geometrical optics (MGO) is a recently developed ray-tracing framework to accurately compute the wavefield behavior near a caustic (turning point or focal point), where traditional ray-tracing breaks down. However, MGO has thus far been restricted to having real-valued wavevectors. This is disadvantageous because often upon crossing a caustic from the ‘illuminated’ region to the ‘shadow’ region, two real-valued rays coalesce into one complex-valued ray corresponding to the transition from propagating to evanescent behavior. One can distinguish caustics as having either ‘illuminated shadows’ or ‘proper shadows’ – the former corresponds to when the shadow still contains real-valued rays (albeit in a fewer quantity than in the illuminated region), while the latter corresponds to when the shadow contains no real-valued rays. Here, by means of examples, we show how MGO can be used to model both types of shadows. First, for illuminated shadows we show that MGO can actually be used ‘as is’, provided a corrected quadrature-angle bias is used compared to that proposed in the original references. This is then implemented and demonstrated in a recently developed MGO ray-tracing code. Second, we show that for proper shadows, the MGO formalism can still be used if the symplectic rotation matrix that removes caustics along rays is allowed to be complex-valued. In both cases, strong agreement is seen between the MGO and the exact solution, demonstrating the potential of MGO for improving the predictive capability of ray-tracing codes and laying the foundations for modeling more complicated evanescent phenomena such as tunnelling with MGO.
@@ Acknowledgements
The work presented here is supported in part by the Carlsberg Foundation, grant CF23-0181. This work has also been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them.